Solve the following system:
3x+2y=-1
5x+4y= 0

Answer by mananth(16946)

3 x + 2 y = 1 .............1

5 x + 4 y = 0 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-6 x -4 y = -2
5 x + 4 y = 0
Add the two equations
-1 x = -2
/ -1
x = 2
plug value of x in (1)
3 x + 2 y = 1
6 + 2 y = 1
2 y = 1 -6
2 y = -5
y = -3
x=2,y=-3